Write an equation in slope intercept form of the line satisfying the given conditions:
The line has an x-intercept at -4 and is parallel to the line containing (3,1) and (2,6).

Answer by mananth(16946)

x1 y1 x2 y2
3 1 2 6

slope m = (y2-y1)/(x2-x1)
( 6 - 1 )/( 2 - 3 )
( 5 / -1 )
m= -5
A Parallel line will have the same slope =-5
Plug value of the slope and point ( -4,0 ) in
Y = m x + b
0.00 = 20 + b
b= 0 - 20
b= -20
So the equation will be
Y = -5x-20


m.ananth@hotmail.ca